OpenStudy (jenniferjuice):

Given the function f(x) = 4(x+3) − 5, solve for the inverse function when x = 3. −21 −16 −10 −1

OpenStudy (anonymous):

solve \[4(x+3)-5=3\]

OpenStudy (jenniferjuice):

4x+12-5=3 4x+12=8 4x=-4 -4/4 x=-1

OpenStudy (jenniferjuice):

oh wait so the final answer is negative one?

zepdrix (zepdrix):

Mmmmm yah that sounds right! Keep on juicin' Jen!!

OpenStudy (jenniferjuice):

lol :D thanks !